# Watermark Anything with Localized Messages

### v1

@@ Abstract
Image watermarking methods are not tailored to handle small watermarked areas. This restricts applications in real-world scenarios where parts of the image may come from different sources or have been edited. We introduce a deep-learning model for localized image watermarking, dubbed the Watermark Anything Model (WAM). The WAM embedder imperceptibly modifies the input image, while the extractor segments the received image into watermarked and non-watermarked areas and recovers one or several hidden messages from the areas found to be watermarked. The models are jointly trained at low resolution and without perceptual constraints, then post-trained for imperceptibility and multiple watermarks. Experiments show that WAM is competitive with state-of-the art methods in terms of imperceptibility and robustness, especially against inpainting and splicing, even on high-resolution images. Moreover, it offers new capabilities: WAM can locate watermarked areas in spliced images and extract distinct 32-bit messages with less than 1 bit error from multiple small regions – no larger than 10% of the image surface – even for small $256\times 256$ images.
[…]
[Code and models]https://github.com/facebookresearch/watermark-anything
@@ 1 Introduction
Invisible image watermarking embeds information into image pixels in a way that is imperceptible to the human eye and yet robust. It was initially developed for intellectual property and copy protection, such as by Hollywood studios for DVDs. However, the applications of watermarking are evolving, particularly in light of the recent development of generative AI models (Kušen and Strembeck, 2018). Regulatory acts such as the White House executive order (USA, 2023), the Californian bill, the EU AI Act (Parliament and Council, 2024), and Chinese AI governance rules (of the People’s Republic of China, 2023) require AI-generated content to be easily identifiable. They all cite watermarking as either compulsory or a recommended measure to detect and label AI-generated images.
@@ 2 Related Work
was first designed for copyright protection (Cox et al., 2007). Traditional methods operate either in the spatial domain (Van Schyndel et al., 1994; Nikolaidis and Pitas, 1998) or in the frequency domain modifying components through an invertible transform (e.g., Fourier, Wavelet) (Cox et al., 1997; Urvoy et al., 2014). These methods were progressively replaced by deep-learning ones which remove the expert need to crafting the transforms such that the embedding is imperceptible and robust. The first approach (Vukotić et al., 2018, 2020; Fernandez et al., 2022; Kishore et al., 2022; Chen et al., 2022) embeds the watermark into the representations of an off-the-shelf pre-trained model. Our method is inspired by the second approach, which jointly trains deep learning-based architectures to embed and extract watermarks while being robust to augmentations seen during training, such as HiDDeN (Zhu et al., 2018) and followers (Zhang et al., 2019b; Luo et al., 2020; Tancik et al., 2020; Ma et al., 2022; Bui et al., 2023b, a; Pan et al., 2024).
@@
Very few papers in the literature deal with watermarking objects in picture (Bas et al., 2001; Barni et al., 2005). This trend was abandoned together with the object-oriented MPEG-4 video codec.
@@
relies on semi-fragile watermarking. Areas where the watermark is not recovered are deemed tampered. This idea appears early in the literature (Kundur and Hatzinakos, 1999; Lin et al., 2000), although deep-learning methods offer significant improvements (Asnani et al., 2022). There is a trade-off between the granularity of the localization of the tampered areas and the semi-fragility of the watermark. The biggest difficulty is to design a watermark robust to benign transformations (e.g., image compression) but fragile to malicious editing.
@@
An extended related work with more details on the methods mentioned above is presented in App. 8.
@@ 3 Detection, Localization, and Message Extraction
DBSCAN offers the advantage of not requiring a pre-defined number of clusters. Instead, we need to specify two parameters: (1) $\text{min}_{\text{samples}}$, the minimum number of points for a cluster to be considered valid, and (2) $\varepsilon$, the maximum distance between two samples for one to be considered as in the neighborhood of the other. For further details on the DBSCAN algorithm, see App. 9.1.
@@ 4.1 The model
WAM considers two joint models: a watermark embedder $\mathrm{emb}_{\theta}$ and a watermark extractor $\mathrm{ext}_{\theta}$, where $\theta$ gathers the parameters of these functions. The embedder defines the watermark procedure for hiding a message into an image, while the extractor detects watermarking and decodes messages. This section discusses our choices for the different pieces that constitute WAM. App. 10.1 gives the technical details of the network architectures.
@@ 4.2 Pre-training models for localized message embedding and extraction
The first step randomly samples a mask among full masks, rectangles, irregular shapes, or object segmentation maps (possibly provided by the training dataset). These masks are inverted with probability $0.5$. It yields a mask $r_{\text{mask}}\in[0,1]^{h\times w}$. Figure 2 shows examples of masks, with more samples displayed in Figure 6 of App. 10.2. The first step then computes the spliced image as $x_{\text{masked}}=r_{\text{mask}}\odot x_{\text{m}}+(1-r_{\text{mask}})\odot x$, where $\odot$ denotes component-wise multiplication.
The second step of the augmentation applies a processing typical from real-world image editing including geometric transformations (identity, resize, crop, rotate, horizontal flip, perspective) and valuemetric adjustments (JPEG, Gaussian blur, median filter, brightness, contrast, saturation, hue). The geometric transformation alters the positions of watermarked pixels, so it is also applied to the mask to keep track of the watermarked areas. This adjusted mask is the ground truth for the localization and it is denoted $y^{\text{det},\star}$ in the sequel. These augmentations are relatively strong to reflect the transformations found on media sharing websites, as opposed to professional photo publications. Additional details and discussions on augmentations are postponed to App. 10.4.
@@ 4.3 Post-training for imperceptibility and multiple watermarks
The Just-Noticeable-Difference (JND) map is a hand-crafted model of the minimum artifact perceivable by a human at every pixel. For instance, artifacts on flat, uniform areas are more visible than on textured ones. The JND map was introduced by Chou and Li (1995) and later used by Wu et al. (2017). App. 9.2 gives more details on its computation.
@@
Since the first training phase (Sec. 4.2) uses a single message, it tends to decode a constant message across pixels even on image regions with different messages. The second training phase introduces several masks to address this limitation. The masks are generated as before with distinct random messages in each of the masked areas. The number of disjoint masks goes from 1 to 3 with probabilities 0.6, 0.2, and 0.2 respectively (see App. 10.2).
@@ 5.1 Implementation details
We provide here further implementation details using the notations introduced in Sec. 4.1, and the rest of the architectures are described in App. 10.1. All experiments are run with $n_{\text{bits}}=32$. There is a total of $1.1$M parameters for the embedder and $96$M for the extractor (equivalent to a ViT-base). In our experiments, we found it necessary to increase the size of the extractor compared to the embedder. This is probably because the extractor has more tasks to handle, such as segmenting and extracting multiple messages, while the embedder only needs to embed one message into an image. Also, it is important for the embedding process to be quick since it happens on the user-side, so we keep it voluntarily small to be fast and usable at scale. We train our model on the MS-COCO training set with blurred faces (Lin et al., 2014), that contains 118,000 images, many of them with segmentation masks. We train at resolution $h\times w=256\times 256$, and with $f=8$ (so $h^{\prime}\times w^{\prime}=32\times 32$). The first training phase (Sec. 4.2) is optimized with AdamW (Kingma, 2014; Loshchilov, 2017) with a linear warmup of the learning rate in $5$ epochs from $1\times 10^{-6}$ to $1\times 10^{-4}$ and a cosine annealing to $1\times 10^{-6}$. We set $\lambda_{\text{dec}}=10$, $\lambda_{\text{det}}=1$, $\alpha=0.3$. We train with a batch size of $16$ per GPU for 300 epochs using 8 V100 GPUs which takes roughly 2 days. The second training phase (Sec. 4.3) further trains the model with the JND attenuation for 200 epochs, hiding up to 3 messages per image using either randomly sampled rectangles or segmentation masks. During this phase, $\alpha_{\text{JND}}=2$. App. 10.2 provides details on the generation of masks used during both training phases. For the extraction of multiple messages, we use the Scikit-learn DBSCAN implementation (Pedregosa et al., 2011).
@@ 5.2 Quality
Table 1 evaluates quantitatively the difference between the watermarked and original images with PSNR, SSIM and LPIPS (Zhang et al., 2018) metrics. We adapt the methods so that they are comparable in terms of imperceptibility. We show qualitative examples for COCO and DIV2k images in App. 12. The perceptual quality is good: the JND map successfully modulates the watermark signal in areas where the eye is not sensitive. Nevertheless, repetitive and regular patterns can be observed in some images when the bright or textured areas are big enough (furs of animals for instance).
@@ 5.3 Detection and decoding
We benchmark WAM’s performance against several watermarking methods: DCTDWT (Al-Haj, 2007), HiDDeN (Zhu et al., 2018), TrustMark (Bui et al., 2023a), SSL (Fernandez et al., 2022), FNNS (Kishore et al., 2022). We also compare WAM to generation-time watermarking methods (Fernandez et al., 2023a; Wen et al., 2023) in App. 11.1. Although this list is not exhaustive, it covers the main types of state-of-the-art methods. For HiDDeN, we use a model which hides 48 bits. FNNS, SSL, and DCTDWT can hide messages or arbitrary length, we choose to hide 48 bits as well. We use the first 16 bits for detection and 32 bits for message decoding. For detection, an image is flagged as watermarked if at most 1 bit is wrongly decoded. This corresponds to a theoretical False Positive Rate (FPR) of $2.6\times 10^{-4}$, i.e., $2.6$ out of 10,000 images are falsely flagged as watermarked on average. TrustMark (Bui et al., 2023a) directly outputs a boolean (watermarked or not). We use the version that hides $40$ bits, but only use the first 32 bits. For all evaluations, we apply the watermark embedding and extraction at the original image resolution as described in the high-resolution paragraph of Sec. 4.1. Finally, for WAM, an image is flagged if $s_{\text{det}}$ (Eq. 1) is higher than 0.07. This value empirically delivers an approximately similar FPR on the COCO validation set.
@@
We evaluate the robustness against various transformations: geometric (flip, crop, perspective, and rotation), valuemetric (adjustments in brightness, hue, contrast or saturation, median or Gaussian filtering, and JPEG compression), splicing: scenarios where only 10% of the image is watermarked superimposed onto the original or an other background. We also evaluate the robustness against inpainting, using the LaMa (Suvorov et al., 2022) model to inpaint areas of the watermarked image specified by random or segmentation masks. For COCO, we use the union of all segmentation masks for each image which in average corresponds to 30% of the image, and for DIV2k we use a randomly generated mask (because there are no segmentation masks) that covers around 35% of the image. We give details on the compared baselines in App. 10.3 and on the transformations in App. 10.4.
@@
2(a) presents the detection and decoding results for low/mid-resolution images, averaged over the first 10k images of the COCO validation set – the detailed results for every augmentation used to form the different groups are in App. 11.3. 2(b) does the same for high-resolution images from the DIV2k (Timofte et al., 2018) validation set. For both distributions, WAM is competitive and even shows improved robustness on classical geometric/valuemetric transformations. Most importantly, WAM performs better on splicing or inpainting where other methods fail to achieve more than 90% TPR and Bit acc. This is true even if WAM was not explicitly trained on high resolution images or to be robust against inpainting. We also show examples of WAM’s detection masks after inpainting in Fig. 10 of App. 11, where some modified parts of images are not detected as watermarked anymore.
@@
Table 2: Detection and decoding after image editing (detailed in Sec. 10.4). We show the bit accuracy (Bit acc.) between the encoded and decoded messages, the proportion of images correctly deemed watermarked (TPR), and the proportion of non-watermarked images falsely detected as watermarked (FPR), in %. Since HiDDeN, DCTDWT, SSL and FNNS do not naturally provide a detection result, we hide 48 bits and reserve 16 bits for detection, and flag an image as watermarked if it has strictly less than two bits incorrectly decoded (these baselines are detailed in Sec. 5.3).
@@ 5.4 Localization
WAM and EditGuard (Zhang et al., 2024) are the only methods that provide watermark localization. We therefore consider images at fixed resolution $512\times 512$ (unlike in Sec. 5.3) to align with the setup of EditGuard. We nevertheless observe similar results at different resolutions for WAM. We focus on the COCO validation set and splice centered watermarked rectangles of various sizes within the images to ensure a well-controlled experiment. We then either apply no transformation or crop the upper left part of the image (25%), which is then resized to the original size. By doing so, the proportion of watermarked pixels is still the same as in the spliced image before cropping, which allows us to evaluate the robustness of the localization. Figure 7 of App. 10.5 illustrates this protocol.
@@ 5.5 Multiple watermarks
We compare the detection and decoding of multiple watermarks before and after the second training phase of WAM (Sec. 4.3). We embed up to five distinct messages into five separate 10% areas of every image (first resized to $256$ to ease the experiments). This is done by feeding the image several times to WAM’s embedder, then pasting the different watermarked areas onto the original image. These areas are squares disposed in a checkerboard pattern, which ensures that the watermarked areas do not overlap and have the same size (see Fig. 8 of App. 10.6). This is an arbitrary choice made to remove confounding factors during evaluation, although the training does not require the watermarked areas to be squared nor to have the same size. Following the methodology detailed in Sec. 3, we apply the DBSCAN algorithm to the vector outputs $\tilde{m}_{i}$ of the extractor corresponding to all pixels $i$ identified as watermarked ($y^{\text{det}}_{i}>\tau$). It identifies clusters corresponding to different watermarked regions without requiring the number of hidden messages to be known in advance. We use $\tau=0.5$ to threshold the watermarked pixels, and we choose a rather strict setup with $\varepsilon=1$ and $\text{min}_{\text{samples}}=1000$ pixels (around $2\%$ of the image) in our evaluations, resulting from an hyperparameter search detailed in Fig. 9 of App. 11.2. This means that a cluster will be considered only if it has at least 2% of the image’s pixels, and that the maximum distance between two predicted messages to be neighbors is 1.
@@
This method remains effective after some image manipulations; extended quantitative results are presented in Fig. 9 of App. 11.2. For example, WAM achieves 85% mIoU even when the images have been horizontally flipped and the contrast adjusted. Furthermore, the bit accuracy, averaged across $5$ messages totaling 160 bits, exceeds 95% under the same conditions. However, WAM fails when JPEG compression is also added on top of these two transformations. Illustrative examples of the identified clusters in various scenarios can be found in Fig. 8 of App. 10.6.

### v2

@@ Abstract
Image watermarking methods are not tailored to handle small watermarked areas. This restricts applications in real-world scenarios where parts of the image may come from different sources or have been edited. We introduce a deep-learning model for localized image watermarking, dubbed the Watermark Anything Model (WAM). The WAM embedder imperceptibly modifies the input image, while the extractor segments the received image into watermarked and non-watermarked areas and recovers one or several hidden messages from the areas found to be watermarked. The models are jointly trained at low resolution and without perceptual constraints, then post-trained for imperceptibility and multiple watermarks. Experiments show that WAM is competitive with state-of-the art methods in terms of imperceptibility and robustness, especially against inpainting and splicing, even on high-resolution images. Moreover, it offers new capabilities: WAM can locate watermarked areas in spliced images and extract distinct 32-bit messages with less than 1 bit error from multiple small regions – no larger than 10% of the image surface – even for small $256\times 256$ images. Training and inference code and model weights are available at github.com/facebookresearch/watermark-anything.
@@ 1 Introduction
Invisible image watermarking embeds information into image pixels in a way that is imperceptible to the human eye and yet robust. It was initially developed for intellectual property and copy protection, such as by Hollywood studios for DVDs. However, the applications of watermarking are evolving, particularly in light of the recent development of generative AI models (Kušen & Strembeck, 2018). Regulatory acts such as the White House executive order (USA, 2023), the Californian bill, the EU AI Act (Parliament & Council, 2024), and Chinese AI governance rules (of the People’s Republic of China, 2023) require AI-generated content to be easily identifiable. They all cite watermarking as either compulsory or a recommended measure to detect and label AI-generated images.
@@ 2 Related Work
was first designed for copyright protection (Cox et al., 2007). Traditional methods operate either in the spatial domain (Van Schyndel et al., 1994; Nikolaidis & Pitas, 1998) or in the frequency domain modifying components through an invertible transform (e.g., Fourier, Wavelet) (Cox et al., 1997; Urvoy et al., 2014). These methods were progressively replaced by deep-learning ones which remove the expert need to crafting the transforms such that the embedding is imperceptible and robust. The first approach (Vukotić et al., 2018; 2020; Fernandez et al., 2022; Kishore et al., 2022; Chen et al., 2022) embeds the watermark into the representations of an off-the-shelf pre-trained model. Our method is inspired by the second approach, which jointly trains deep learning-based architectures to embed and extract watermarks while being robust to augmentations seen during training, such as HiDDeN (Zhu et al., 2018) and followers (Zhang et al., 2019b; Luo et al., 2020; Tancik et al., 2020; Ma et al., 2022; Bui et al., 2023b; a; Pan et al., 2024).
@@
Very few papers in the literature deal with watermarking objects in pictures (Bas et al., 2001; Barni et al., 2005). This trend was abandoned together with the object-oriented MPEG-4 video codec.
@@
relies on semi-fragile watermarking. Areas where the watermark is not recovered are deemed tampered. This idea appears early in the literature (Kundur & Hatzinakos, 1999; Lin et al., 2000), although deep-learning methods offer significant improvements (Asnani et al., 2022). There is a trade-off between the granularity of the localization of the tampered areas and the semi-fragility of the watermark. The biggest difficulty is to design a watermark robust to benign transformations (e.g., image compression) but fragile to malicious editing.
@@
An extended related work with more details on the methods mentioned above is presented in App. B.
@@ 3 Detection, Localization, and Message Extraction
DBSCAN offers the advantage of not requiring a pre-defined number of clusters. Instead, we need to specify two parameters: (1) $\text{min}_{\text{samples}}$, the minimum number of points for a cluster to be considered valid, and (2) $\varepsilon$, the maximum distance between two samples for one to be considered as in the neighborhood of the other. For further details on the DBSCAN algorithm, see App. C.1.
@@ 4.1 The model
WAM considers two joint models: a watermark embedder $\mathrm{emb}_{\theta}$ and a watermark extractor $\mathrm{ext}_{\theta}$, where $\theta$ gathers the parameters of these functions. The embedder defines the watermark procedure for hiding a message into an image, while the extractor detects watermarking and decodes messages. This section discusses our choices for the different pieces that constitute WAM. App. D.1 gives the technical details of the network architectures.
@@ 4.2 Pre-training models for localized message embedding and extraction
The first step randomly samples a mask among full masks, rectangles, irregular shapes, or object segmentation maps (possibly provided by the training dataset). These masks are inverted with probability $0.5$. It yields a mask $r_{\text{mask}}\in[0,1]^{h\times w}$. Figure 2 shows examples of masks, with more samples displayed in Figure 6 of App. D.2. The first step then computes the spliced image as $x_{\text{masked}}=r_{\text{mask}}\odot x_{\text{m}}+(1-r_{\text{mask}})\odot x$, where $\odot$ denotes component-wise multiplication.
The second step of the augmentation applies a processing typical from real-world image editing including geometric transformations (identity, resize, crop, rotate, horizontal flip, perspective) and valuemetric adjustments (JPEG, Gaussian blur, median filter, brightness, contrast, saturation, hue). The geometric transformation alters the positions of watermarked pixels, so it is also applied to the mask to keep track of the watermarked areas. This adjusted mask is the ground truth for the localization and it is denoted $y^{\text{det},\star}$ in the sequel. These augmentations are relatively strong to reflect the transformations found on media sharing websites, as opposed to professional photo publications. Additional details and discussions on augmentations are postponed to App. D.4.
@@ 4.3 Post-training for imperceptibility and multiple watermarks
The Just-Noticeable-Difference (JND) map is a hand-crafted model of the minimum artifact perceivable by a human at every pixel. For instance, artifacts on flat, uniform areas are more visible than on textured ones. The JND map was introduced by Chou & Li (1995) and later used by Wu et al. (2017). App. C.2 gives more details on its computation.
@@
Since the first training phase (Sec. 4.2) uses a single message, it tends to decode a constant message across pixels even on image regions with different messages. The second training phase introduces several masks to address this limitation. The masks are generated as before with distinct random messages in each of the masked areas. The number of disjoint masks goes from 1 to 3 with probabilities 0.6, 0.2, and 0.2 respectively (see App. D.2).
@@ 5.1 Implementation details
We provide here further implementation details using the notations introduced in Sec. 4.1, and the rest of the architectures are described in App. D.1. All experiments are run with $n_{\text{bits}}=32$. There is a total of $1.1$M parameters for the embedder and $96$M for the extractor (equivalent to a ViT-base). In our experiments, we found it necessary to increase the size of the extractor compared to the embedder. This is probably because the extractor has more tasks to handle, such as segmenting and extracting multiple messages, while the embedder only needs to embed one message into an image. Also, it is important for the embedding process to be quick since it happens on the user-side, so we keep it voluntarily small to be fast and usable at scale. We train our model on the MS-COCO training set with blurred faces (Lin et al., 2014), that contains 118,000 images, many of them with segmentation masks. We train at resolution $h\times w=256\times 256$, and with $f=8$ (so $h^{\prime}\times w^{\prime}=32\times 32$). The first training phase (Sec. 4.2) is optimized with AdamW (Kingma, 2014; Loshchilov, 2017) with a linear warmup of the learning rate in $5$ epochs from $1\times 10^{-6}$ to $1\times 10^{-4}$ and a cosine annealing to $1\times 10^{-6}$. We set $\lambda_{\text{dec}}=10$, $\lambda_{\text{det}}=1$, $\alpha=0.3$. We train with a batch size of $16$ per GPU for 300 epochs using 8 V100 GPUs which takes roughly 2 days. The second training phase (Sec. 4.3) further trains the model with the JND attenuation for 200 epochs, hiding up to 3 messages per image using either randomly sampled rectangles or segmentation masks. During this phase, $\alpha_{\text{JND}}=2$. App. D.2 provides details on the generation of masks used during both training phases. For the extraction of multiple messages, we use the Scikit-learn DBSCAN implementation (Pedregosa et al., 2011).
@@ 5.2 Quality
Table 1 evaluates quantitatively the difference between the watermarked and original images with PSNR, SSIM and LPIPS (Zhang et al., 2018) metrics. We adapt the methods so that they are comparable in terms of imperceptibility. We show qualitative examples for COCO and DIV2k images in App. F. The perceptual quality is good: the JND map successfully modulates the watermark signal in areas where the eye is not sensitive. Nevertheless, repetitive and regular patterns can be observed in some images when the bright or textured areas are big enough (furs of animals for instance).
@@ 5.3 Detection and decoding
We benchmark WAM’s performance against several watermarking methods: DCTDWT (Al-Haj, 2007), HiDDeN (Zhu et al., 2018), TrustMark (Bui et al., 2023a), SSL (Fernandez et al., 2022), FNNS (Kishore et al., 2022). We also compare WAM to generation-time watermarking methods (Fernandez et al., 2023a; Wen et al., 2023) in App. E.1. Although this list is not exhaustive, it covers the main types of state-of-the-art methods. For HiDDeN, we use a model which hides 48 bits. FNNS, SSL, and DCTDWT can hide messages or arbitrary length, we choose to hide 48 bits as well. We use the first 16 bits for detection and 32 bits for message decoding. For detection, an image is flagged as watermarked if at most 1 bit is wrongly decoded. This corresponds to a theoretical False Positive Rate (FPR) of $2.6\times 10^{-4}$, i.e., $2.6$ out of 10,000 images are falsely flagged as watermarked on average. TrustMark (Bui et al., 2023a) directly outputs a boolean (watermarked or not). We use the version that hides $40$ bits, but only use the first 32 bits. For all evaluations, we apply the watermark embedding and extraction at the original image resolution as described in the high-resolution paragraph of Sec. 4.1. Finally, for WAM, an image is flagged if $s_{\text{det}}$ (Eq. 1) is higher than 0.07. This value empirically delivers an approximately similar FPR on the COCO validation set.
@@
We evaluate the robustness against various transformations: geometric (flip, crop, perspective, and rotation), valuemetric (adjustments in brightness, hue, contrast or saturation, median or Gaussian filtering, and JPEG compression), splicing: scenarios where only 10% of the image is watermarked superimposed onto the original or an other background. We also evaluate the robustness against inpainting, using the LaMa (Suvorov et al., 2022) model to inpaint areas of the watermarked image specified by random or segmentation masks. For COCO, we use the union of all segmentation masks for each image which in average corresponds to 30% of the image, and for DIV2k we use a randomly generated mask (because there are no segmentation masks) that covers around 35% of the image. We give details on the compared baselines in App. D.3 and on the transformations in App. D.4.
@@
2(a) presents the detection and decoding results for low/mid-resolution images, averaged over the first 10k images of the COCO validation set – the detailed results for every augmentation used to form the different groups are in App. E.3. 2(b) does the same for high-resolution images from the DIV2k (Timofte et al., 2018) validation set. For both distributions, WAM is competitive and even shows improved robustness on classical geometric/valuemetric transformations. Most importantly, WAM performs better on splicing or inpainting where other methods fail to achieve more than 90% TPR and Bit acc. This is true even if WAM was not explicitly trained on high resolution images or to be robust against inpainting. We also show examples of WAM’s detection masks after inpainting in Fig. 10 of App. E, where some modified parts of images are not detected as watermarked anymore.
@@
Table 2: Detection and decoding after image editing (detailed in Sec. D.4). We show the bit accuracy (Bit acc.) between the encoded and decoded messages, the proportion of images correctly deemed watermarked (TPR), and the proportion of non-watermarked images falsely detected as watermarked (FPR), in %. Since HiDDeN, DCTDWT, SSL and FNNS do not naturally provide a detection result, we hide 48 bits and reserve 16 bits for detection, and flag an image as watermarked if it has strictly less than two bits incorrectly decoded (these baselines are detailed in Sec. 5.3).
@@ 5.4 Localization
WAM and EditGuard (Zhang et al., 2024) are the only methods that provide watermark localization. We therefore consider images at fixed resolution $512\times 512$ (unlike in Sec. 5.3) to align with the setup of EditGuard. We nevertheless observe similar results at different resolutions for WAM. We focus on the COCO validation set and splice centered watermarked rectangles of various sizes within the images to ensure a well-controlled experiment. We then either apply no transformation or crop the upper left part of the image (25%), which is then resized to the original size. By doing so, the proportion of watermarked pixels is still the same as in the spliced image before cropping, which allows us to evaluate the robustness of the localization. Figure 7 of App. D.5 illustrates this protocol.
@@ 5.5 Multiple watermarks
We compare the detection and decoding of multiple watermarks before and after the second training phase of WAM (Sec. 4.3). We embed up to five distinct messages into five separate 10% areas of every image (first resized to $256$ to ease the experiments). This is done by feeding the image several times to WAM’s embedder, then pasting the different watermarked areas onto the original image. These areas are squares disposed in a checkerboard pattern, which ensures that the watermarked areas do not overlap and have the same size (see Fig. 8 of App. D.6). This is an arbitrary choice made to remove confounding factors during evaluation, although the training does not require the watermarked areas to be squared nor to have the same size. Following the methodology detailed in Sec. 3, we apply the DBSCAN algorithm to the vector outputs $\tilde{m}_{i}$ of the extractor corresponding to all pixels $i$ identified as watermarked ($y^{\text{det}}_{i}>\tau$). It identifies clusters corresponding to different watermarked regions without requiring the number of hidden messages to be known in advance. We use $\tau=0.5$ to threshold the watermarked pixels, and we choose a rather strict setup with $\varepsilon=1$ and $\text{min}_{\text{samples}}=1000$ pixels (around $2\%$ of the image) in our evaluations, resulting from an hyperparameter search detailed in Fig. 9 of App. E.2. This means that a cluster will be considered only if it has at least 2% of the image’s pixels, and that the maximum distance between two predicted messages to be neighbors is 1.
@@
This method remains effective after some image manipulations; extended quantitative results are presented in Fig. 9 of App. E.2. For example, WAM achieves 85% mIoU even when the images have been horizontally flipped and the contrast adjusted. Furthermore, the bit accuracy, averaged across $5$ messages totaling 160 bits, exceeds 95% under the same conditions. However, WAM fails when JPEG compression is also added on top of these two transformations. Illustrative examples of the identified clusters in various scenarios can be found in Fig. 8 of App. D.6.